OpenStudy (christos):

Linear Algebra, say I have a matrix to which I want to perform gauss-Jordan elimination. I see 2 identical rows. Can I substract one from the other at the same time and get 2 zero rows ?

ILovePuppiesLol (ilovepuppieslol):

@Mehek14 @Mehek14 @Mehek14

ILovePuppiesLol (ilovepuppieslol):

idk this sorry :(

rishavraj (rishavraj):

if u mean \[\left[\begin{matrix}a & b & c \\ d & e & f \\ d & e & f\end{matrix}\right]\] after subtraction becomes \[\left[\begin{matrix}a & b & c \\ 0 & 0 & 0 \\ 0 & 0 & 0\end{matrix}\right]\] u cant do tht

OpenStudy (christos):

:/ ok